OpenStudy (anonymous):

@jim_thompson5910 Graph the circle with center at (-2, 1) and diameter 18. Label the center and at least four points on the circle. Write the equation of the circle.

OpenStudy (anonymous):

equation of the circle with centre (a,b) and radius r is \[(x-a)^2+(y-b)^2 =r^2\]

jimthompson5910 (jim_thompson5910):

if the diameter is 18, then the radius is 9 (since 9 is half of 18)

jimthompson5910 (jim_thompson5910):

so you graph this by plotting the center (-2,1) Then going 9 units in every direction from the center to graph the circle

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

your equation would be \[\Large (x-h)^2+(y-k)^2 =r^2\] \[\Large (x-(-2))^2+(y-1)^2 =9^2\] \[\Large (x+2)^2+(y-1)^2 =81\] after plugging in the center and radius